OpenStudy (anonymous):

simplify completely x^2-4x+4/x^2+10x+25 multiplied by x+5/x^2+3x-10

OpenStudy (anonymous):

\[\Large \frac{x^2-4x+4}{x^2+10x+25} \cdot \frac{x+5}{x^2+3x-10}\] factorize .. and you'll get : \[\Large \frac{(x-2)^2}{(x+5)^2} \cdot \frac{x+5}{(x-2)(x+5)}\] now simplify ...

OpenStudy (anonymous):

\[\frac{x-2}{(x+5)^2}\]

OpenStudy (anonymous):

(with the restriction that \[x \neq 2\]